# Simultaneous FUV and NUV observations of T Tauri stars with UVIT/AstroSat: probing accretion process in young stars

### v1

@@ 3.3 Relation between H luminosity and accretion luminosity
In this study, we have determined excess UV flux by fitting the excess emission with the black-body spectrum. The flux enclosed within the fitted black-body spectrum is nothing but the excess UV emission due to accretion in CTTS and chromospheric activity in the WTTS. Hence, the bolometric luminosity of the black-body spectrum provides the direct measure of accretion luminosity in CTTS. We used this luminosity value to calculate mass accretion rate in CTTS. Mass accretion rate (()) and accretion luminosity are related by
@@
where M and R are stellar mass and radius and R is the disk truncation radius from which the gas falls onto the star. R is typically assumed to be 5 R (Gullbring et al., 1998). The error introduced by this assumption on the measured mass accretion rates, considering that R for a pre-main sequence star can span from 3 to 8 R, is less than 20%. Therefore, the above equation can be written as
@@
We have considered R/M 5 R /M and used above equation 3 to estimate . The values are listed in the last column of Table 3. We have only calculated the mass accretion rate for the stars classified as CTTS. L and estimated from excess UV flux provides a direct measure of these parameters.
@@
EW(H) is often used as a proxy for accretion (White & Basri, 2003). However, EW(H) is secondary indicator for accretion. Therefore, we tried to find the correlation between L from UV excess and . We used extinction corrected Pan-STARRS r band flux density as a proxy for the continuum flux density underlying H line and the distance obtained from Gaia-DR3 to calculate from average EW(H) values listed in Table 2. The continuum flux density at H is given as (Mathew et al., 2018),
@@
where and r is the extinction corrected Pan-STARRS magnitude. The extinction is obtained from the SED fitting as listed in Table 3. In Figure 6, we plotted L estimated from UV excess vs that estimated from , which shows a linear correlation (red line) between them with a slope of 1.210.17 and intercept of 2.180.56. There are also previously reported correlations between luminosity of H (L) and accretion luminosity (L) (Vogt et al., 1994; Herczeg & Hillenbrand, 2008; Dahm, 2008; Fang et al., 2009; Ingleby et al., 2013). We compare our relation with that estimated by previous studies in Table 4 and noticed that our relation matches quite well with the slope and intercept values with the previous estimation from the literature. However, we get a relatively large error in its intercept value, which could be due to the small sample size. A large sample with simultaneous observations in UV and optical spectroscopy is required for a better correlation. We have also used this relation to estimate L from L for all the CTTS listed in Table 2. This help us to calculate L for the TTS with no FUV observations for which we could not estimate UV luminosity. We then also calculated using Equation 3. The values of L and are listed in Table 2. We notice that J0414, CIDA 1, V836 Tau, FN Tau, and BS Tau are the slowly accreting CTTSs with a mass accretion rate a few times 10 Myr, while FM Tau, CW Tau, and FO Tau are found to have comparatively higher accretion rate (a few times 10 Myr).
@@ 3.4 Relation between UV excess and IR excess
The NIR color excess if taken as evidence for the presence of circumstellar disks around TTS. The strength of the infrared colour excess indicates the amount of dust and gas in these disks. A strong excess suggests a substantial amount of material in the disk, while a weaker excess may indicate a less massive or more evolved disk. The strong excess in FUV and NUV suggests the presence of significant accretion onto the stars, while the high FUV and relatively low or no NUV excess suggest that the UV emission originates from chromospheric activity. FUV emission from the star plays a significant role in heating gas disks and driving massive thermal winds which deplete the disk material. Hence, excess in (FUV NUV) color in WTTS stars will tell us if the stars are dominant in chromospheric activity, while for CTTS, it tells us whether the stars are dominant in FUV emission. Also, the comparison between UV color and IR color helps us to identify the changes in disk morphology (primordial gas-rich disks or evolved or transitioning to a debris disk) from FUV bright stars to FUV faint stars. The UV colour index is plotted against J-K, JH and HK infrared color excess from 2MASS observations in Figure 7. HK excess is considered to be a good estimator of the accretion rate (Meyer et al., 1997). The colour excess is defined as: )], in magnitude unit. For CTTS, we notice that stars with lower IR colour excess, i.e., with weaker inner disks, appear bluer in the (FUV NUV) color, which means they have stronger FUV emission. This suggests that strong FUV emission might have caused the depletion of the gas and dust in the circumstellar disk in CTTS. The WTTS seem to be following a different slope relation compared to CTTS but with two WTTS data points, it is not so conclusive. A larger sample of CTTS and WTTS is required for better understanding of this relation.
@@ 4 Summary
In this study, we present the accretion properties of eleven TTSs in the TMC, namely FM Tau, BS Tau, V836 Tau, HD283782, V773 Tau, CW Tau, FN Tau, FO Tau, Anon 1, CIDA 1, and 2MASS J04141188+2811535. This is the first UVIT study of T Tauri stars and highlights the significance of simultaneous multi-band UV observations of TTSs to study their accretion properties. Out of 11, seven are detected in both FUV and NUV bands, while the remaining four are detected only in NUV. We report, for the first time, the UV photometry of 2MASS J04141188+2811535 (not detected in GALEX, which could be due to small exposure time), and the FUV photometry of FN Tau.
@@
We noticed there are two categories of TTSs based on the excess UV emission: one shows strong excess emission (10 in NUV and 10 FUV) over the photosphere; the other with excess emission mostly in FUV (10) and a little excess in NUV (10). We classify the TTSs with strong excess emission in both FUV and NUV as CTTS, and the other group as WTTS. This study shows that UV excess can be used as a tool to distinguish CTTS and WTTS.
We found that classification of TTS as classical or weak-line based on the spectral type vs EW(H) relation (Barrado y Navascués & Martín, 2003; White & Basri, 2003) matches well with that made based on UV excess. Our classification also matches well with the literature. From our analysis, we reconfirm that V836 Tau as CTTS and for the first time we report BS Tau and J0414 as CTTS, and Anon 1 as WTTS.
We found that H luminosity and accretion luminosity (from UV luminosity) are linearly correlated with a slope of 1.210.17 and an intercept of 2.180.56. The correlation also matches well with previous estimates by (Herczeg & Hillenbrand, 2008; Dahm, 2008; Fang et al., 2009; Manara et al., 2012). From the mass accretion rate calculation based on UV luminosity and H luminosity, we found that CW Tau, FM Tau and FO Tau are going through a strong accretion phase with an accretion rate Myr, while other CTTS are accreting at a slower rate ( a few times 10 Myr). The study brings out the importance of multi-wavelength SED analysis and simultaneous FUV and NUV observation to get a better estimation of the accretion luminosity and accretion rate of TTS.
Comparing the UV colour with the IR colour, we notice that stars with higher IR colour excess, i.e., the presence of a substantial amount of material in the disk, appear redder in the (FUV NUV) colour, and vice versa. This suggests that stronger FUV emission might have caused the depletion of the gas in the disk. However, a study with a large number sample is required for a better understanding of this relation.

### v2

@@ 3.3 Relation between H luminosity and accretion luminosity
In this study, we have determined excess UV flux by fitting the excess emission with the black-body spectrum. The flux enclosed within the fitted black-body spectrum is nothing but the excess UV emission due to accretion in a CTTS and chromospheric activity in a WTTS. Hence, the bolometric luminosity of the black-body spectrum provides the direct measure of accretion luminosity in a CTTS. We used this luminosity value to calculate the mass accretion rates in CTTSs. Mass accretion rate (()) and accretion luminosity are related by the following equation by Gullbring et al. (1998):
@@
where M∗ and R∗ are stellar mass and radius and Rin is the disk truncation radius from which the gas falls onto the star. Rin is typically assumed to be 5 R∗ (Gullbring et al., 1998). The error introduced by this assumption on the measured mass accretion rates, considering that Rin for a pre-main sequence star can span from 3 to 8 R∗, is less than 20%. Therefore, the above equation can be written as
@@
We have considered R∗/M∗ 5 R⊙ /M⊙ and used above equation 3 to estimate . The values are listed in the last column of Table 3. We have only calculated the mass accretion rate for the stars classified as CTTSs. Lacc and estimated from excess UV flux provides a direct measure of these parameters.
@@
EW(H) is often used as a proxy for accretion (White & Basri, 2003). However, EW(H) is a secondary indicator for accretion. Therefore, we tried to find the correlation between Lacc from UV excess and . We used extinction-corrected Pan-STARRS r band flux density as a proxy for the continuum flux density underlying the Hα line and the distance obtained from Gaia-DR3 to calculate from average EW(Hα) values listed in Table 2. The continuum flux density at H is given as (Mathew et al., 2018),
@@
where and r0 is the extinction corrected Pan-STARRS magnitude. The extinction is obtained from the SED fitting as listed in Table 3. In Figure 6, we plotted Lacc estimated from UV excess vs that estimated from , which shows a linear correlation (red line) between them with a slope of 1.200.22 and intercept of 2.160.70. The uncertainties associated to shown in Figure 6 mainly signifies the range in values from . The uncertainties in Lacc or UV luminosities are obtained from the observational errors in flux and in distance estimations, not from the errors in best-fitted model parameters. As both observed flux errors and distance errors are very small, the uncertainties in Lacc are also found to be very small. There are also previously reported correlations between L and Lacc (Vogt et al., 1994; Herczeg & Hillenbrand, 2008; Dahm, 2008; Fang et al., 2009; Ingleby et al., 2013). We compare our relation with that estimated by previous studies in Table 4 and noticed that our relation matches quite well with the slope and intercept values with the previous estimation from the literature. However, we get a relatively large error in its intercept value, which could be due to the small sample size and non-simultaneous observation of and UV. A large sample with simultaneous observations in UV and optical spectroscopy is required for a better correlation. We have also used this relation to estimate Lacc from L for all the CTTSs listed. This helps us to calculate Lacc for the TTSs with no FUV observations for which we could not estimate UV luminosity. We then also calculated using Equation 3. The values of Lacc and are listed in columns 6 and 7 of Table 2, respectively. The uncertainties associated with Lacc and indicate the range in these values due to the range in . We notice that J0414, CIDA 1, V836 Tau, and BS Tau are the slowly accreting CTTSs with a mass accretion rate a few times 10-9 M⊙yr-1, while FM Tau, CW Tau, and FO Tau are found to have comparatively higher accretion rate (a few times 10-8 M⊙yr-1).
@@ 3.4 Relation between UV excess and IR excess
The NIR color excess is taken as evidence for the presence of circumstellar disks around TTSs. The strength of the infrared color excess indicates the amount of dust and gas in these disks. A strong excess suggests a substantial amount of material in the disk, while a weaker excess may indicate a less massive or more evolved disk. The strong excess in FUV and NUV suggests the presence of significant accretion onto the stars. In contrast, the high FUV and relatively low or no NUV excess suggest that the UV emission originates from chromospheric activity. FUV emission from the star plays a significant role in heating gas disks and driving massive thermal winds which deplete the disk material. Hence, excess in (FUV NUV) color in WTTSs will tell us if the stars are dominant in chromospheric activity, while for CTTSs, it tells us whether the stars are dominant in FUV emission. Also, the comparison between UV color and NIR color helps us to identify the changes in disk morphology (primordial gas-rich disks or evolved or transitioning to a debris disk) from FUV bright stars to FUV faint stars. The UV color index is plotted against JK infrared color excess from 2MASS observations in Figure 7. JK excess is considered to be a good estimator of the accretion rate (Meyer et al., 1997). The color excess is defined as: )], in magnitude unit. For CTTS, we notice that stars with lower NIR color excess, i.e., with weaker inner disks, appear bluer in the (FUV NUV) color, which means they have stronger FUV emission. This suggests that strong FUV emission might have caused the depletion of the gas and dust in the circumstellar disk in CTTSs. The WTTSs seem to be following a different slope relation compared to CTTSs but with two WTTSs data points, it is not so conclusive. A larger sample of CTTSs and WTTSs is required to better understand this relation.
@@ 4 Summary
In this study, we present the accretion properties of 10 TTSs in the TMC, namely FM Tau, BS Tau, V836 Tau, HD283782, V773 Tau, CW Tau, FO Tau, Anon 1, CIDA 1, and 2MASS J04141188+2811535. This is the first UVIT study of T Tauri stars and highlights the significance of simultaneous multi-band UV observations of TTSs to study their accretion properties. Six of 10 are detected in both FUV and NUV bands, while the remaining four are detected only in NUV. We report, for the first time, the UV photometry of 2MASS J04141188+2811535 (not detected in GALEX, which could be due to short exposure time).
@@
We noticed there are two categories of TTSs based on the excess UV emission: one shows strong excess emission (103 in NUV and 107 FUV) over the photosphere; the other with excess emission mostly in FUV (105) and a little excess in NUV (102). We classify the TTSs with strong excess emission in both FUV and NUV as CTTS, and the other group as WTTS. This study shows that UV excess can be used as a tool to distinguish CTTS and WTTS.
We found that the classification of TTS as classical or weak-line based on the spectral type vs EW(H) relation (Barrado y Navascués & Martín, 2003; White & Basri, 2003) matches well with that made based on UV excess. Our classification also matches well with the literature. From our analysis, we reconfirm that V836 Tau is CTTS. For the first time, we report that BS Tau, CIDA 1 and 2MASS J04141188+2811535 are still actively accreting and classify them as CTTS, and Anon 1 as non-accreting or WTTS.
We found that H luminosity and accretion luminosity (from UV luminosity) are linearly correlated with a slope of 1.200.22 and an intercept of 2.160.70. The correlation also matches well with previous estimates by (Herczeg & Hillenbrand, 2008; Dahm, 2008; Fang et al., 2009; Manara et al., 2012). From the mass accretion rate calculation based on UV luminosity and H luminosity, we found that CW Tau, FM Tau and FO Tau are going through a strong accretion phase with an accretion rate M⊙yr-1, while other CTTS are accreting at a slower rate ( a few times 10-9 M⊙yr-1). The study brings out the importance of multi-wavelength SED analysis and simultaneous FUV and NUV observation to get a better estimation of the accretion luminosity and accretion rate of TTS.
Comparing the UV color with the NIR color, we notice that stars with higher NIR color excess, i.e., the presence of a substantial amount of material in the disk, appear redder in the (FUV NUV) color, and vice versa. This suggests that stronger FUV emission might have caused the depletion of the gas in the disk. However, a study with a large number sample is required for a better understanding of this relation.